OpenStudy (anonymous):

a sample of copper wire with a mass of 1.659 g is heated with excess powdered sulfur until the copper has combined with the sulfur, and all the extra sulfur has been burned off. The final product (a sulfide of copper ) has a mass of 2.080 g. Calculate the empirical formula of the product, and name it.

OpenStudy (jebonna):

First lets find out the mass of sulfur used: Copper = 1.659g Sulfur = 2.080 - 1.659 = 0.421g Now we can have to convert the mass of each substance into the moles of each substance using this equation: mass/molar mass Copper: 1.659g/63.5g mol-1 = 0.0261mol Sulfur: 0.421g/32.1g mol-1 = 0.0131mol Now what we have to do is divide the one with the smallest moles by each answer we got, so 0.0131 is the smaller amount of moles, (this gives us how many of each element is in the compound) so first we do: Sulfur = 0.0131/0.0131 = 1 And now we divide the smallest amount of moles by our other answer (which is the moles of copper) Copper = 0.0261/0.0131 = 2 (rounded up) From this, we have found out that there is 1 sulfur in the compound and 2 copper. Therefore the product would look like this: Cu2S I hope this helps!

OpenStudy (anonymous):

thank you jebonna

OpenStudy (jebonna):

No problem :)